OpenStudy (anonymous):

how do i find out if the ordered pair is a solution of an ordered pair? 2x+2y is less than or equal to 0; (-1,-1), (1,1)

OpenStudy (anonymous):

solution of an inequality..

OpenStudy (anonymous):

Put -1=x and -1=y to check if (-1,-1) is a solution. Similarly, put x=1 and y=1 to check if (1,1) is a solution. 2(-1)+2(-1)=-2-2=-4 which is less than 0 so (-1,-1) is a solution. Also, 2(1)+2(1)=2+2=4 which is NOT less than 0 so (1,1) is NOT a solution.

OpenStudy (anonymous):

okay, thank you, that seems simple enough!

OpenStudy (anonymous):

It's a pleasure to be of any help :D